OpenStudy (aravindg):

Note:This is not a question !! Tutorial 2 Topic:Maths tricks (contd) hope you like it .Do give ur feedback!

OpenStudy (aravindg):

Square of any 2 digit number __________________________ Let me explain this trick by taking examples 67^2 = [6^2][7^2]+20*6*7 = 3649+840 = 4489 similarly 25^2 = [2^2][5^2]+20*2*5 = 425+200 = 625 Take one more example 97^2 = [9^2][7^2]+20*9*7 = 8149+1260 = 9409 Here [] is not an operation, it is only a separation between initial 2 and last 2 digits Here an extra case arises Consider the following examples for that 91^2 = [9^2][1^2]+20*9*1 = 8101+180 = 8281 nice isnt it ?

OpenStudy (anonymous):

yeah !!! but is it applicable for every N ?

OpenStudy (aravindg):

holds for N upto 100 ...

OpenStudy (anonymous):

N- natural numbers?

OpenStudy (shubhamsrg):

any good reason ? if you can give ?

OpenStudy (shubhamsrg):

well i may add,,hope you dont mind (afterall its your tutorial) we have (10 a + b)^2 = 100a^2 + b^2 + 20.a.b a and b both are single digits,,max value b^2 can take is 81.. 2 0's are there in the end when 100 gets multiplied by a^2 and hence is the trick..

OpenStudy (anonymous):

WOW

OpenStudy (anonymous):

thats brillliant

OpenStudy (goformit100):

@AravindG jahapana tussi great ho TOFA KUBUL KARO.... :)

OpenStudy (lgbasallote):

hmm haven't seen these things for months....

OpenStudy (swag):

awesome

OpenStudy (anonymous):

well, i find this trick much easier, to find the square of any 2-digit no. suppose 10x+y 1. add the unit's digit to the no., ie. 10x+y+y 2.multiply the result with the ten's digit as:- , (10x+2y)*(10x) 3. add the square of the unit's digit to the result obtained after doing the 2nd step. and you'll get the desired square. eg. \[43^{2}= (43+3)(40)+9=1849\]

OpenStudy (anonymous):

btw @AravindG your tricks are awesome :)

OpenStudy (aravindg):

@shreya. thanks :) ...that trick is new to me and its interesting ! Thanks for sharing !

OpenStudy (anonymous):

:))))))